Suppose a basketball player has made 231 out of 361 free throws. If the player makes the next 2 free throws, I will pay you $31. Otherwise you pay me $21. Step 1 of 2 : Find the expected value of the proposition. Round your answer to two decimal places. Losses must be expressed as negative values.
Answer #1

P(winning) = (231/361)*(231/361) = 0.40946

P(loss) = 1-0.40946 = 0.59054

expected value of the proposition = (31*0.40946)+(-21*0.59054) = 0.29 (ans)